# SportsGPT: An LLM-driven Framework for Interpretable Sports Motion Assessment and Training Guidance

### v1

@@ Abstract
Existing intelligent sports analysis systems mainly focus on "scoring and visualization," often lacking automatic performance diagnosis and interpretable training guidance. Recent advances of Large Language Models (LMMs) and motion analysis techniques provide new opportunities to address the above limitations. In this paper, we propose SportsGPT, an LLM-driven framework for interpretable sports motion assessment and training guidance, which establishes a closed loop from motion time-series input to professional training guidance. First, given a set of high-quality target models, we introduce MotionDTW, a two-stage time series alignment algorithm designed for accurate keyframe extraction from skeleton-based motion sequences. Subsequently, we design a Knowledge-based Interpretable Sports Motion Assessment Model (KISMAM) to obtain a set of interpretable assessment metrics (e.g., insufficient extension) by constrasting the keyframes with the targe models. Finally, we propose SportsRAG, a RAG-based training guidance model based on Qwen3. Leveraging a 6B-token knowledge base, it prompts the LLM to generate professional training guidance by retrieving domain-specific QA pairs. Experimental results demonstrate that MotionDTW significantly outperforms traditional methods with lower temporal error and higher IoU scores. Furthermore, ablation studies validate the KISMAM and SportsRAG, confirming that SportsGPT surpasses general LLMs in diagnostic accuracy and professionalism.
@@ 1 Introduction
Despite notable progress, current "intelligent" sports applications still face several critical bottlenecks. First, most systems remain limited to basic "scoring + visualization" functionalities and simple digitization pipelines [3, 5, 10, 11, 12, 13, 14]. Although these tools present kinematic and kinetic data efficiently, they lack automated diagnosis, reliable interpretation, personalized training guidance. A simple scoring or scaling function is insufficient for athletes to identify the occurrence of incorrect motions or to understand how to rectify them. Moreover, many amateur athletes and coaches often lack the scientific expertise required to design systematic rehabilitation strategies. Traditional models and visualization tools fail to provide personalized and practical training guidance to improve the athelte’s performance. Finally, low automation in critical stages-specifically keyframe extraction-severely bottlenecks system efficiency and precision. While keyframe extraction is essential for bridging raw video and biomechanical analysis [15, 16, 17, 18], existing works often rely on time-consuming manual annotation. Meanwhile, classic algorithms like Dynamic Time Warping (DTW) [19, 20, 21] struggle with inter-individual variability and speed disparities, resulting in poor robustness and a continued dependence on manual intervention.
@@
Extensive experiments validate the efficacy of our framework. The proposed MotionDTW achieves state-of-the-art alignment precision with average temporal errors of 0.91 and 1.59 frames on the 100m sprint (Aceleration phase) and standing long jump, respectively, while maintaining superior computational efficiency and stability than baselines. In downstream assessment tasks, our method yields diagnostic results with significantly higher Intersection over Union (IoU) with ground truth compare to standard DTW-based approaches. Furthermore, human evaluations demonstrate that SportsGPT surpasses general-purpose LLMs in accuracy, comprehensiveness, professionalism, and feasibility, highlighting its practical value for professional sports training.
In summary, the main contributions of this paper can be summarized as follows:
@@ 3.1 Motion Evaluation module
To enhance robustness against local noise and capture short-term action trends, we apply a sliding window mechanism. For time $t$, features from the preceding and succeeding $w$ frames ($w=4$) are concatenated to form the temporal window feature. Enriching contextual information and reducing single-frame jitter effects.
@@
Two-Stage Matching Strategy. To handle redundant actions in input videos, we propose a two-stage matching strategy that first performs subsequence serach and then conducts keyframe alignment.By restricting the search space to a band with radius $r$ around the diagonal, FastDTW [21] reduces the time complexity from $O(N^{2})$ to $O(N)$, enabling real-time processing of high-dimensional feature sequences.
@@
A kinematic metric anomaly ($D^{k}>0$) may indicate one or multiple technical issues. We define a set $\mathcal{P}={P_{1},\dots,P_{n}}$ of $n$ common problems, from "nappropriate start step length" to "center of mass too high at start." A mapping matrix $W\in\mathbb{R}^{K\times M}$ specifies the contribution of each metrics anomaly to each problem. For example, "flight time larger than reference" may correspond to $P_{1}$ or $P_{2}$, while "support leg tibia angle larger than reference" may indicate $P_{9}$ or $P_{11}$.
@@ 3.2 SportsRAG: RAG-based training guidance model
Although KISMAM effectively maps kinematic deviations into structured assessment metrics (e.g., insufficient start extension), these numerical indicators lack contextual explanation and actionable guidance for end-users. Direct use of general Large Language Models (LLMs) often leads to domain knowledge gaps and severe hallucinations. To address this, we propose SportsRAG, a RAG-based training guidance model based on Qwen3-8B. Unlike traditional fine-tuning approaches, SportsRAG leverages a massive external knowledge base to ground generation in verified expert data, ensuring scientific rigor and minimizing hallucinations.
To serve as the retrieval source for RAG, we constructed a 6B-token knowledge base, categorized into three levels of granularity to ensure coverage and depth:
  • Theoretical Foundation (Academic Corpus): We curated over 200 authoritative textbooks (e.g., Sports Biomechanics, Exercise Physiology) and core sports science journals from the past decade, providing the theoretical bounds for generation.
  • Practical Experience (Expert QA): To align with real-world needs, we aggregated common inquiries from expert daily training and rehabilitation routines, alongside queries from web users. Through rigorous manual annotation, we constructed a dataset of 50,000 high-quality QA pairs, serving as precise retrieval anchors for specific scenarios.
  • Reference Standards (Historical Reports): We incorporated 1,000 professional analysis reports historically generated by student athletes using the AI system. These records exemplify the complete mapping from quantitative assessment data to specific training schemes and guidance, providing proven templates for reliable report generation.
@@ 4.1 Performance on Training Guidance
To rigorously assess utility, we conducted a double-blind evaluation where three senior specialists scored 60 randomly sampled reports on a 1-5 Likert scale across four dimensions: Accuracy (interpretation correctness), Comprehensiveness (inference depth), Professionalism (expert-level specificity), and Feasibility (actionability of quantitative details).
@@
4.2 Performance on Motion Assessment
To assess the practical value of precise keyframe extraction, we integrated our method into the downstream KISMAM assessment pipeline. We utlize the acceleration phase test set to extract keyframes via FastDTW (Baseline) and MotionDTW (Ours), which were then input into KISMAM to generate the Top-6 core biomechanical diagnosis. We evaluate performance using the Intersection over Union (IoU) metric. Let $\mathcal{D}$ denote the set of biomechanical diagnostic rules. Given the model’s Top-6 prediction set $\mathcal{S}_{pred}\subset\mathcal{D}$ and the expert ground truth $\mathcal{S}_{gt}\subset\mathcal{D}$, the IoU is defined as shown in Eq. (6):
@@ 4.3 Ablation Study
To validate the robustness of our alignment algorithm, we calculated the Average Keyframe Error across all five test scenarios (four phases of the 100m sprint and the standing long jump). We systematically analyzed the contribution of each module and the impact of the search strategy. The quantitative results are summarized in Table 2.
@@
  • w/o Dynamic Features (Ang. Vel/Acc): The error nearly doubles to 2.74 frames. This confirms that in high-speed sports, static joint angles alone are insufficient, and higher-order derivatives are critical for capturing rapid temporal changes.
  • w/o Joint Weights: Removing joint weights increases the error to 2.61 frames, indicating that focusing on core kinematic chains (e.g., lower limbs for sprinting) effectively suppresses noise from irrelevant body parts.
  • w/o Multi-Joint: Relying on single-joint analysis results in a moderate error increase to 1.79 frames, suggesting that holistic body coordination is beneficial for robust representation.
@@ 5 Conclusion and Future Work
In this paper, we presented an LLM-driven framework for interpretable sports motion assessment and training guidance that bridges the gap between quantitative biomechanical characteristics and actionable expert advice. By synergizing MotionDTW for precise spatiotemporal alignment, KISMAM for semantic diagnostic mapping, and a RAG-Based training guidance model, our system effectively mitigates the interpretability gap and hallucination issues prevalent in existing solutions. Empirical results confirm that our approach achieves state-of-the-art alignment precision and generates professional-grade training prescriptions comparable to those of human experts, offering a robust paradigm for intelligent coaching.

### v2

@@ Abstract
Existing intelligent sports analysis systems mainly focus on "scoring and visualization," often lacking automatic performance diagnosis and interpretable training guidance. Recent advances in Large Language Models (LLMs) and motion analysis techniques provide new opportunities to address the above limitations. In this paper, we propose SportsGPT, an LLM-driven framework for interpretable sports motion assessment and training guidance, which establishes a closed loop from motion time-series input to professional training guidance. First, given a set of high-quality target models, we introduce MotionDTW, a two-stage time series alignment algorithm designed for accurate keyframe extraction from skeleton-based motion sequences. Subsequently, we design a Knowledge-based Interpretable Sports Motion Assessment Model (KISMAM) to obtain a set of interpretable assessment metrics (e.g., insufficient extension) by contrasting the keyframes with the target models. Finally, we propose SportsRAG, a RAG-based training guidance model built upon Qwen3. Leveraging a 6B-token knowledge base, it prompts the LLM to generate professional training guidance by retrieving domain-specific QA pairs. Experimental results demonstrate that MotionDTW significantly outperforms traditional methods with lower temporal error and higher IoU scores. Furthermore, ablation studies validate the KISMAM and SportsRAG, confirming that SportsGPT surpasses general LLMs in diagnostic accuracy and professionalism.
@@ 1 Introduction
Despite notable progress, current "intelligent" sports applications still face several critical bottlenecks. First, most systems remain limited to basic "scoring + visualization" functionalities and simple digitization pipelines [3, 5, 10, 11, 12, 13, 14]. Although these tools present kinematic and kinetic data efficiently, they lack automated diagnosis, reliable interpretation, and personalized training guidance. A simple scoring or scaling function is insufficient for athletes to identify the occurrence of incorrect motions or to understand how to rectify them. Moreover, many amateur athletes and coaches often lack the scientific expertise required to design systematic rehabilitation strategies. Traditional models and visualization tools fail to provide personalized and practical training guidance to improve the athlete’s performance. Finally, low automation in critical stages-specifically keyframe extraction-severely bottlenecks system efficiency and precision. While keyframe extraction is essential for bridging raw video and biomechanical analysis [15, 16, 17, 18], existing works often rely on time-consuming manual annotation. Meanwhile, classic algorithms like Dynamic Time Warping (DTW) [19, 20, 21] struggle with inter-individual variability and speed disparities, resulting in poor robustness and a continued dependence on manual intervention.
@@
Extensive experiments validate the efficacy of our framework. The proposed MotionDTW achieves state-of-the-art alignment precision with average temporal errors of 0.91 and 1.59 frames on the 100m sprint (Acceleration phase) and standing long jump, respectively, while maintaining superior computational efficiency and stability to baselines. In downstream assessment tasks, our method yields diagnostic results with significantly higher Intersection over Union (IoU) with ground truth compare to standard DTW-based approaches. Furthermore, human evaluations demonstrate that SportsGPT surpasses general-purpose LLMs in accuracy, comprehensiveness, professionalism, and feasibility, highlighting its practical value for professional sports training.
The main contributions of this paper are summarized as follows:
@@ 3.1 Motion Evaluation module
To enhance robustness against local noise and capture short-term action trends, we apply a sliding window mechanism. For time $t$, features from the preceding and succeeding $w$ frames ($w=4$) are concatenated to form the temporal window feature. This enriches contextual information and reduces single-frame jitter effects.
@@
Two-Stage Matching Strategy. To handle redundant actions in input videos, we propose a two-stage matching strategy that first performs subsequence search and then conducts keyframe alignment.By restricting the search space to a band with radius $r$ around the diagonal, FastDTW [21] reduces the time complexity from $O(N^{2})$ to $O(N)$, enabling real-time processing of high-dimensional feature sequences.
@@
A kinematic metric anomaly ($D^{k}>0$) may indicate one or multiple technical issues. We define a set $\mathcal{P}={P_{1},\dots,P_{n}}$ of $n$ common problems, from "Inappropriate start step length" to "center of mass too high at start." A mapping matrix $W\in\mathbb{R}^{K\times M}$ specifies the contribution of each metrics anomaly to each problem. For example, "flight time larger than reference" may correspond to $P_{1}$ or $P_{2}$, while "support leg tibia angle larger than reference" may indicate $P_{9}$ or $P_{11}$.
@@ 3.2 SportsRAG: RAG-based training guidance model
Although KISMAM effectively maps kinematic deviations into structured assessment metrics (e.g., insufficient start extension), these numerical indicators lack contextual explanation and actionable guidance for end-users. Direct use of general Large Language Models (LLMs) often leads to domain knowledge gaps and severe hallucinations. To address this, we propose SportsRAG, a RAG-based training guidance model built upon Qwen3-8B. Unlike traditional fine-tuning approaches, SportsRAG leverages a massive external knowledge base to ground generation in verified expert data, ensuring scientific rigor and minimizing hallucinations.
To serve as the retrieval source for RAG, we construct a 6B-token knowledge base, categorized into three levels of granularity to ensure coverage and depth:
  • Theoretical Foundation (Academic Corpus): We curate over 200 authoritative textbooks (e.g., Sports Biomechanics, Exercise Physiology) and core sports science journals from the past decade, providing the theoretical bounds for generation.
  • Practical Experience (Expert QA): To align with real-world needs, we aggregate common inquiries from expert daily training and rehabilitation routines, alongside queries from web users. Through rigorous manual annotation, we constructed a dataset of 50,000 high-quality QA pairs, serving as precise retrieval anchors for specific scenarios.
  • Reference Standards (Historical Reports): We incorporate 1,000 professional analysis reports historically generated by student athletes using the AI system. These records exemplify the complete mapping from quantitative assessment data to specific training schemes and guidance, providing proven templates for reliable report generation.
@@ 4 Experiments
4.1 Performance on Motion Assessment
To assess the practical value of precise keyframe extraction, we integrate our method into the downstream KISMAM assessment pipeline. We utilize the acceleration phase test set to extract keyframes via FastDTW (Baseline) and MotionDTW (Ours), which are then input into KISMAM to generate the Top-6 core biomechanical diagnosis. We evaluate performance using the Intersection over Union (IoU) metric. Let $\mathcal{D}$ denote the set of biomechanical diagnostic rules. Given the model’s Top-6 prediction set $\mathcal{S}_{pred}\subset\mathcal{D}$ and the expert ground truth $\mathcal{S}_{gt}\subset\mathcal{D}$, the IoU is defined as shown in Eq. (6):
@@ 4.2 Performance on Training Guidance
To rigorously assess utility, we conduct a double-blind evaluation where three senior specialists score 60 randomly sampled reports on a 1-5 Likert scale across four dimensions: Accuracy (interpretation correctness), Comprehensiveness (inference depth), Professionalism (expert-level specificity), and Feasibility (actionability of quantitative details).
@@ 4.3 Ablation Study
To validate the robustness of our alignment algorithm, we calculate the Average Keyframe Error across all five test scenarios (four phases of the 100m sprint and the standing long jump). We systematically analyze the contribution of each module and the impact of the search strategy. The quantitative results are summarized in Table 2.
@@
Removing dynamic features (angular velocity and acceleration) causes the error to nearly double to 2.74 frames, confirming that static joint angles alone are insufficient for high-speed sports and that higher-order derivatives are essential for capturing rapid temporal changes. Similarly, eliminating joint weights increases the error to 2.61 frames, indicating that prioritizing core kinematic chains-such as the lower limbs in sprinting-effectively suppresses noise from irrelevant body parts. Furthermore, relying solely on single-joint analysis results in a moderate error increase to 1.79 frames, suggesting that holistic body coordination is beneficial for robust representation.
@@ 5 Conclusion and Future Work
In this paper, we present an LLM-driven framework for interpretable sports motion assessment and training guidance that bridges the gap between quantitative biomechanical characteristics and actionable expert advice. By synergizing MotionDTW for precise spatiotemporal alignment, KISMAM for semantic diagnostic mapping, and a RAG-Based training guidance model, our system effectively mitigates the interpretability gap and hallucination issues prevalent in existing solutions. Empirical results confirm that our approach achieves state-of-the-art alignment precision and generates professional-grade training prescriptions comparable to those of human experts, offering a robust paradigm for intelligent coaching.